OpenStudy (anonymous):

Solve: 2(x + 1) = −4

OpenStudy (ankit042):

can you expand the left hand side?

HanAkoSolo (jamierox4ev3r):

you would use the property of distribution 2x+2=-4

HanAkoSolo (jamierox4ev3r):

then, here is the rest of the process 2x+2=-4 -2 -2 2x=-6 /2 /2 x=-3

HanAkoSolo (jamierox4ev3r):

darn @FilthyMcNasty u beat me :S

HanAkoSolo (jamierox4ev3r):

but back to the original point: x=-3 do you understand @Sav_1012 ?